OpenStudy (anonymous):

Given right triangle A w/a hypotenuse length of 3y + x and a leg of y-x, and right triangle B, w/ a hypotense length of y+5 and a leg length of x+5, for what values of x and y are the triangles congruent by HL?

OpenStudy (anonymous):

Congruent by HL ryt? So that means that their hypothenuse and the leg are equal... So.. 3y+x = y+5 y-x = x+5 from here we get... y = 2x+5 Substitute; 3(2x+5)+x = 2x +5+5 7x+15 = 2x+10 5x = -5 x= -1 Substitute: y =2(-1)+5 = 3